2x+3y=3
3x-7y= -30
solve system by addition method. Solve for x,y

Answer by Alan3354(69443)

2x+3y=3
3x-7y= -30
solve system by addition method.
============
Multiply the 1st eqn by 3, and the 2nd by -2.
then add.
---
6x + 9y = 9 --- 1st eqn times 3
-6x+14y = 60 --- 2nd eqn times -2
--------------------------------------------- Add
23y = 69
y = 3
------------------
2x + 3y = 3
2x + 9 = 3
2x = -6
x = -3